# Exploring Simple-Population and Multiple-Population Globular Clusters in the Outer Galactic Halo using the Hubble Space Telescope

### v1

@@ Abstract
The pseudo two-color diagram, known as chromosome map (ChM), is a valuable tool for identifying globular clusters (GCs) that consist of single or multiple stellar populations (MPs). Recent surveys of Galactic GCs using the ChM have provided stringent observational constraints on the formation of GCs and their stellar populations. However, these surveys have primarily focused on GCs at moderate distances from the Galactic center and composed of MPs.
In this paper, we present the first detailed study of the stellar composition of four GCs in the outer halo of the Milky Way: Arp 2, Ruprecht 106, Terzan 7, and Terzan 8. Our analysis is based on high-precision photometry obtained from images collected with the Hubble Space Telescope in the F275W, F336W, F438W, F606W, and F814W bands. We find that Ruprecht 106 and Terzan 7 are composed solely of a single stellar population, whereas Arp 2 and Terzan 8 host both first- and second-population stars. In these clusters, the second population comprises about half and one-third of the total number of GC stars, respectively.
The results from this paper and the literature suggest that the threshold in the initial GC mass, if present, should be smaller than approximately . The first-population stars of Arp 2 and Terzan 8, along with the stars of the simple-population GCs Ruprecht 106 and Terzan 7, exhibit intrinsic color spreads corresponding to [Fe/H] variations of approximately 0.05 – 0.30 dex. This indicates that star-to-star metallicity variations are a common feature of star clusters, regardless of the presence of MPs.
@@ 1 Introduction
In this context, the identification of simple-population (SP) GCs, namely clusters with chemically homogeneous stars, casts doubt on the universal occurrence of MPs, thus challenging the view of MPs as a phenomenon inherent to the proto-cluster environment. The best prototypes of SP GCs are Ruprecht 106 and Terzan 7. For the first cluster, the spectroscopic analysis of nine red giant stars by Villanova et al. (2013) revealed homogeneous light-element abundance. Moreover, the analysis of the cluster HB morphology by Milone et al. (2014), and the photometry in the F336W, F438W, and F814W bands by Dotter et al. (2018), suggested that the color distribution of RGB stars was consistent with that of a single stellar population. Similarly, evidence based on ground-based photometry that the star cluster Terzan 7 contains a single stellar population has been provided by Lagioia et al. (2019b).
@@
As for the metallicity, the selected clusters span a wide range of iron content, ranging from the metal poor cluster Terzan 8 ( dex; Carretta et al., 2014) to the sub-solar metallicity cluster Terzan 7 ( dex Tautvaišienė et al., 2004). We notice that the average metallicity values reported in the (Harris, 1996, 2010 update) catalog span a similar interval, with [Fe/H] dex for Terzan 8 and [Fe/H] dex for Terzan 7.
All the targets are relatively small clusters with present-day mass M , except Terzan 8, which is the most massive GC with M (Baumgardt & Hilker, 2018, 2023 update). Baumgardt and collaborators have also provided an estimate of the initial mass of the target clusters 111In 2018, Baumgardt & Hilker provided advanced estimates for the initial masses of the Galactic GCs. Their determinations face several uncertainties which depend on our limited understanding of the evolutionary history of the Milky Way, and therefore of its tidal field. Other uncertainties are related to the knowledge of the primordial mass segregation in clusters and their early chemo-dynamical evolution, in particular the amount of mass lost through interaction with the molecular clouds from which they originated. We notice that the aforementioned GC mass estimates are obtained by assuming a time-independent Galactic gravitational potential and no initial mass segregation., , that for three out of four clusters is smaller (albeit marginally in the case of Ruprecht 106) than , a threshold value proposed by Milone et al. (2020) for the onset of MPs. It is worth mentioning that according to studies based on the analysis of the CN and CH indices in low resolution spectra of relatively small and distant globulars, the initial mass threshold for the onset of MP might be lower than the value suggested by Milone et al. (2020) and approximately equal to (Salinas & Strader, 2015; Tang et al., 2021; Huang et al., 2024).
Finally, concerning the last criterion, several studies indicate a probable ex-situ genesis of Arp 2, Terzan 7 and Terzan 8, which are associated to the Sgr dSph (Law & Majewski, 2010; Bellazzini et al., 2003) and likely still bound to it (Bellazzini et al., 2020; Shirazi et al., 2023).
@@ 3 The UV-optical-NIR CMDs
Figure 2 displays the versus pseudo-CMDs of our targets, arranged as in Fig. 1. The pseudo-CMDs of the two GCs in the top row show the largest color spread among the four targets. In particular, we observe a distinct split along the entire RGB extension in Arp 2, whereas the RGB stars of Terzan 8 with similar magnitudes display a broad color spread, with indications of bifurcation.
In both the cases, the pseudo-color broadening of RGB stars is taken as evidence that these clusters harbor MPs. A spurious effect due to random errors in the pseudo-color is ruled out by the extension of the color error bars. On the other side, the RGB spread visible in the CMDs of Ruprecht 106 and Terzan 7 is comparable with the error in color along the whole RGB extension. This suggests that Arp 2 and Terzan 8 may harbor MPs with different light-element abundances, while Ruprecht 106 and Terzan 7 may not.
@@ 4 The chromosome maps
To further investigate the stellar populations along the RGB of the studied GCs, we constructed a pseudo color-color diagram called ChM. In this diagram, stars with different content of light elements such as helium, carbon, nitrogen, oxygen, and magnesium, occupy specific regions (Milone et al., 2017; Marino et al., 2019b). As a result, a ChM gives the possibility to tag distinct stellar populations in a GC with a precision comparable with that provided by high-resolution spectroscopy (Marino et al., 2019a). In the subsequent section, we briefly describe the method used to derive the ChMs of the studied GCs, whereas in section 4.2 we presents the analysis of the ChMs and results on the fraction of 1P and 2P stars in the target clusters.
@@ 4.1 Constructing the chromosome maps
In a nutshell, to derive the ChM abscissa we take advantage of the versus CMD, and of the versus pseudo-CMD to derive the ChM ordinate. In both these diagrams, we measure the (pseudo-) color displacement of every RGB star relative to the red boundary of the RGB sequence, and divide it by the broadening of the RGB sequence. The RGB broadening is defined as the (pseudo-) color difference between the red and the blue RGB boundaries at the same F814W star’s magnitude. This result of this process, called verticalization, is a color ratio that is finally multiplied by a factor called intrinsic RGB width.
To derive the intrinsic RGB width, we first measure the broadening of the RGB, or observed RGB width, at a reference F814W magnitude set by convention at 333The F814W luminosity of the MSTO has been determined by using the naive estimator method (Silverman, 1986). First, we selected a magnitude interval in the versus CMD which includes MS and sub-giant branch (SGB) stars, and divided it into a given number of magnitude bins. Then, we computed the median color and magnitude of the stars in each bin. The whole procedure has been repeated -times, with assuming a different value for each cluster ( = 5, 6, 8), keeping the bin-width unchanged but shifting the starting point of the first bin by an amount equivalent to 1/ of the bin-width. For each bin, the resulting median points have been linearly interpolated across the entire initial magnitude interval thus obtaining a raw fiducial line, which has been finally smoothed by boxcar averaging three adjacent points. Finally, the magnitude of the bluest point of the smoothed fiducial line has been taken as the cluster MS turn-off luminosity in F814W band. A similar approach has been adopted for the determination of the RGB boundaries. The blue and red RGB envelopes correspond, respectively, to the fiducial lines interpolating the and percentile of the RGB (pseudo-) color distribution in each bin, across the whole magnitude range covered by the cluster RGB members. Their relative distance of the distance at two F814W magnitudes above the MSTO correspond to the observed RGB width..
Then, since the observed RGB width encompasses both the intrinsic RGB width and the observational error, it is necessary to estimate the contribution of the latter to the observed RGB width. To do that, we simulated the versus and versus diagrams for a SP composed of 10,000 stars. The errors associated to each magnitude of the simulated stars are extracted from a Gaussian distribution with a dispersion corresponding to the median magnitude error of the RGB stars. Finally, the intrinsic RGB width in the color and in the pseudo-color , indicated respectively as and , have been obtained by subtracting in quadrature the RGB widths of the simulated CMD from observed RGB width ones.
@@ 4.2 Multiple stellar populations in the chromosome map
Similarly to Fig. 4, the red dashed line in the ChM of Terzan 8 and Arp 2 separates bona-fide 1P and 2P stars. As far as Ruprecht 106 and Terzan 7 are concerned, our procedure results in a homogeneous composition of 1P stars only.Both Terzan 8 and Arp 2 show a relatively high fraction of 1P stars compared to most of the studied Galactic GCs. In particular, in Terzan 8 of the cluster RGB stars belong to the 1P, thus making it one of the Galactic GCs with the largest fractions of 1P stars. In the case of Arp 2, 1P stars make up a fraction slightly larger than 50% of the RGB cluster stars.
@@ 5 Metallicity variations among first-population stars
In particular, we note that, although Ruprecht 106 does not show any vertical spread, so no presence of 2P stars, the horizontal extension of its ChM covers a interval of mag, significantly larger than that of the photometric errors. Such a phenomenon is even more pronounced in the case of Terzan 8, where the extension of the 1P candidate stars is mag, while in the case of Arp 2 1P and 2P stars seem to span a comparable interval along the x-axis.
Recent works have shown that the horizontal width of 1P stars in the ChMs is a widespread phenomenon among Galactic GCs and demonstrated that 1P stars are not chemically homogeneous (Milone et al., 2015; D’Antona et al., 2016; Milone et al., 2018b). Papers based on both spectroscopy and photometry have shown that the range of 1P stars is mostly due to metallicity variations (e.g. Marino et al., 2019a; Legnardi et al., 2022; Marino et al., 2023).
@@ 6 Summary and conclusions
We derived the versus ChM to search for MPs with different light-element abundances. We find that Arp 2 and Terzan 8 host MPs, whereas Ruprecht 106 and Terzan 7 are SP clusters. This is the first evidence of simple stellar population GCs based on the ChM, and corroborates previous conclusions based on spectroscopy of nine stars of Ruprecht 106 (Villanova et al., 2013), photometry in the U, B, and I bands of RGB stars in both Ruprecht 106 and Terzan 7 (Dotter et al., 2018; Lagioia et al., 2019a), and the HB morphology of Ruprecht 106 (Milone et al., 2014).
@@
It is widely accepted that the pseudo-color width of the RGB correlates with the average cluster metallicity. In GCs with MPs, the RGB width mostly depends on the maximum internal variation in the abundances of certain light elements. This color index is primarily sensitive to the nitrogen content of the stellar populations and, to a lesser extent, to variations in helium, carbon, and oxygen (e.g. Milone et al., 2015; Lagioia et al., 2019a; Jang et al., 2021). In addition to light-element spread, star-to-star metallicity variations can also contribute to the RGB width (e.g. Marino et al., 2019b). We find that the observed RGB widths of Arp 2 and Terzan 8 follow the same trend as the bulk of Galactic GCs in the versus [Fe/H] diagram. On the contrary, the RGB width of Ruprecht 106 and Terzan 7 are much smaller than those of GCs with similar [Fe/H] as expected for clusters with homogeneous abundance of light-elements.
@@
Previous works have suggested that a mass threshold could govern the occurrence of MPs in GCs. Based on analysis of the ChMs in 58 Galactic GCs, Milone et al. (2020) noticed that the MP GCs have masses larger than . Arp 2 is the Galactic GC with the smallest initial mass ( ) for which MPs have been detected through the ChMs. The discovery of MPs in this cluster suggests that the proposed mass threshold, if present, should be smaller than . In this context, it is worth noticing that various SP Magellanic Cloud GCs have initial masses of 1 – 3 , which are larger than those of Arp 2 and Terzan 8. This fact could indicate that, in addition to cluster mass, the GC formation environment, affects the occurrence of MPs with metal-poor GCs having a lower initial-mass threshold for the formation of 2P stars (Krause et al., 2016; Huang et al., 2024). In this regard, a possible interpretation for the non occurrence of MPs in some ex-situ GCs is that the formation environment was not suitable for the formation of different stellar populations. In this sense, as recently pointed out by Huang et al. (2024), GCs formed in dwarf galaxies during later starbursts, therefore relatively younger, can be more metal-rich and less compact, two conditions that disfavor the intra-cluster retention of gas expelled by rapidly evolved 1P stars, and so the birth of 2P stars. As a consequence, the GC formation environment would be a discriminant factors for the onset of MPs (Krause et al., 2016).
An alternative explanation for the lack of MPs is that this phenomenon would only occur in stars less massive than only (Bastian & Lardo, 2018). In such a case case, the reason why Milone et al. (2020) we did not detect 2P stars in some Magellanic Cloud star clusters is that the studies of MPs in these Magellanic Cloud clusters younger than Gyr are exclusively based on RGB stars.
@@
The 1P stars of the multiple-population GCs Arp 2 and Terzan 8 exhibit intrinsic color spreads, which are indicative of star-to-star [Fe/H] variations of and 0.3 dex, respectively. A recent study of 55 Galactic GCs with MPs by Legnardi et al. (2022) has shown that the 1P stars of virtually all clusters exhibit inhomogeneous metallicities, with [Fe/H] abundance variation ranging from less than 0.05 dex to dex. Hence, the results on Arp 2 and Terzan 8 corroborate the evidence that metallicity variations among 1P stars are common features of GCs.
@@
We discovered that the stars of the SP GCs Ruprecht 106 and Terzan 7 exhibit metallicity variations, in close analogy with what is observed in 1P stars of MP GCs.This fact, together with the findings of star-to-star [Fe/H] variations in the open cluster NGC 6791 and in the Gyr old star cluster NGC 1783 (Legnardi et al., 2024), demonstrates that the metallicity variations are not a peculiarity of multiple-population GCs but are also present in SP GCs, open clusters, and intermediate-age massive Magellanic-cloud clusters. We confirm that the maximum iron variation among the 1P stars mildly correlate with cluster mass and metallicity.

### v2

@@ Abstract
The pseudo two-color diagram, known as chromosome map (ChM), is a valuable tool for identifying globular clusters (GCs) that consist of single or multiple stellar populations (MPs). Recent surveys of Galactic GCs using the ChM have provided stringent observational constraints on the formation of GCs and their stellar populations. However, these surveys have primarily focused on GCs at moderate distances from the Galactic center and composed of MPs. In this paper, we present the first detailed study of the stellar composition of four GCs in the outer halo of the Milky Way: Arp 2, Ruprecht 106, Terzan 7, and Terzan 8. Our analysis is based on high-precision photometry obtained from images collected with the Hubble Space Telescope in the F275W, F336W, F438W, F606W, and F814W bands. We find that Ruprecht 106 and Terzan 7 are composed solely of a single stellar population, whereas Arp 2 and Terzan 8 host both first- and second-population stars. In these clusters, the second population comprises about half and one-third of the total number of GC stars, respectively. The results from this paper and the literature suggest that the threshold in the initial GC mass, if present, should be smaller than approximately . The first-population stars of Arp 2 and Terzan 8, along with the stars of the simple-population GCs Ruprecht 106 and Terzan 7, exhibit intrinsic color spreads, likely indicative of [Fe/H] variations of approximately 0.05 – 0.30 dex. This suggests that star-to-star metallicity variations are a common feature of star clusters, regardless of the presence of MPs.
@@ 1 Introduction
In this context, the identification of simple-population (SP) GCs, namely clusters with chemically homogeneous stars, casts doubt on the universal occurrence of MPs, thus challenging the view of MPs as a phenomenon inherent to the proto-cluster environment. The best prototypes of SP GCs are Ruprecht 106 and Terzan 7. For the first cluster, the spectroscopic analyses by Villanova et al. (2013); François (2014) and Frelijj et al. (2021) revealed homogeneous light-element abundance. Moreover, the analysis of the cluster HB morphology by Milone et al. (2014), and the photometry in the F336W, F438W, and F814W bands by Dotter et al. (2018), suggested that the color distribution of RGB stars was consistent with that of a single stellar population. Similarly, evidence based on ground-based photometry that the star cluster Terzan 7 contains a single stellar population has been provided by Lagioia et al. (2019b).
@@
As for the metallicity, the selected clusters span a wide range of iron content, ranging from the metal poor cluster Terzan 8 ( dex; Mottini et al., 2008; Carretta et al., 2014) to the sub-solar metallicity cluster Terzan 7 ( dex Tautvaišienė et al., 2004). We notice that the average metallicity values reported in the (Harris, 1996, 2010 update) catalog span a similar interval, with [Fe/H] dex for Terzan 8 and [Fe/H] dex for Terzan 7.
All the targets are relatively small clusters with present-day mass M , except Terzan 8, which is the most massive GC among them, with M (Baumgardt & Hilker, 2018, 2023 update). Baumgardt and collaborators have also provided an estimate of the initial mass of the target clusters 111In 2018, Baumgardt & Hilker provided advanced estimates for the initial masses of the Galactic GCs. Their determinations face several uncertainties which depend on our limited understanding of the evolutionary history of the Milky Way, and therefore of its tidal field. Other uncertainties are related to the knowledge of the primordial mass segregation in clusters and their early chemo-dynamical evolution, in particular the amount of mass lost through interaction with the molecular clouds from which they originated. We notice that the aforementioned GC initial mass estimates are obtained by assuming a time-independent Galactic gravitational potential and no initial mass segregation., , that for three out of four clusters is smaller (albeit marginally in the case of Ruprecht 106) than , a threshold value proposed by Milone et al. (2020) for the onset of MPs. It is worth mentioning that according to studies based on the analysis of the CN and CH indices in low resolution spectra of relatively small and distant globulars, the initial mass threshold for the onset of MP might be lower than the value suggested by Milone et al. (2020) and approximately equal to (Salinas & Strader, 2015; Tang et al., 2021; Huang et al., 2024).
Finally, concerning the last criterion, several studies indicate a probable ex-situ genesis of Arp 2, Terzan 7 and Terzan 8, which are associated to the Sgr dSph (Bellazzini et al., 2003; Law & Majewski, 2010; Carretta et al., 2014) and likely still bound to its main body (Bellazzini et al., 2020). The origin of Ruprecht 106 is doubtful: while previous studies rule out a link between this cluster and Sgr Pritzl et al. (e.g. 2005), recent kinematic studies instead assess its association to the Sgr stream (Shirazi et al., 2023). Interestingly, Sbordone et al. (2007) showed that the chemical composition of Ruprecht 106 is similar to that of the Sgr field stars. More recently Villanova et al. (2013); François (2014) and Lucertini et al. (2023) have suggested an extra-galactic origin for Ruprecht 106 based on the relative chemical abundance of , r- and s-process elements in the analyzed stars, compared to those typical of GC and field stars of similar metallicity.
@@ 3 The UV-optical-NIR CMDs
Figure 2 displays the versus pseudo-CMDs of our targets, arranged as in Fig. 1. The pseudo-CMDs of the two GCs in the top row show the largest color spread among the four targets. In particular, we observe a distinct split along the entire RGB extension in Arp 2, whereas the RGB stars of Terzan 8 with comparable magnitudes display a broad color spread, with indications of a split.
In both the cases, the pseudo-color broadening of RGB stars is taken as evidence that these clusters harbor MPs. A spurious effect due to random errors in the pseudo-color is ruled out by the size of the color error bars. On the other side, the RGB spread visible in the CMDs of Ruprecht 106 and Terzan 7 is comparable with the error in color along the whole RGB extension. This suggests that Arp 2 and Terzan 8 harbor MPs with different light-element abundances, while Ruprecht 106 and Terzan 7 may not.
@@ 4 The chromosome maps
To further investigate the stellar populations along the RGB of the studied GCs, we constructed a pseudo color-color diagram called ChM. In this diagram, stars with different content of light elements such as helium, carbon, nitrogen, oxygen, and magnesium, occupy specific regions (Milone et al., 2017; Marino et al., 2019b). As a result, a ChM gives the possibility to classify large numbers of stars into various populations and with an accuracy competing with that provided by high-resolution spectroscopy (Marino et al., 2019a). In the subsequent section, we briefly describe the method used to derive the ChMs of the studied GCs, whereas in section 4.2 we presents the analysis of the ChMs and results on the fraction of 1P and 2P stars in the target clusters.
@@ 4.1 Constructing the chromosome maps
In a nutshell, to derive the ChM abscissa we take advantage of the versus CMD, and of the versus pseudo-CMD to derive the ChM ordinate. In both these diagrams, we measure the (pseudo-) color displacement of every RGB star relative to the red boundary of the RGB sequence, and divide it by the broadening of the RGB sequence. The RGB broadening is defined as the (pseudo-) color difference between the red and the blue RGB boundaries at the same F814W star’s magnitude. This result of this process, called verticalization, is a normalized color ratio that is finally multiplied by a factor called intrinsic RGB width.
To derive the intrinsic RGB width, we first measure the broadening of the RGB, or observed RGB width, at a reference F814W magnitude set by convention at 333The F814W luminosity of the MSTO has been determined by using the naive estimator method (Silverman, 1986). First, we selected a magnitude interval in the versus CMD which includes MS and sub-giant branch (SGB) stars, and divided it into a given number of magnitude bins. Then, we computed the median color and magnitude of the stars in each bin. The whole procedure has been repeated -times, with assuming a different value for each cluster ( = 5, 6, 8), keeping the bin-width unchanged but shifting the starting point of the first bin by an amount equivalent to 1/ of the bin-width. For each bin, the resulting median points have been linearly interpolated across the entire initial magnitude interval thus obtaining a raw fiducial line, which has been finally smoothed by boxcar averaging three adjacent points. Finally, the magnitude of the bluest point of the smoothed fiducial line has been taken as the cluster MS turn-off luminosity in F814W band. A similar approach has been adopted for the determination of the RGB boundaries. The blue and red RGB envelopes correspond, respectively, to the fiducial lines interpolating the and percentile of the RGB (pseudo-) color distribution in each bin, across the whole magnitude range covered by the cluster RGB members. Their relative distance of the distance at two F814W magnitudes above the MSTO correspond to the observed RGB width.. Then, since the observed RGB width encompasses both the intrinsic RGB width and the observational error, it is necessary to estimate the contribution of the latter to the observed RGB width. To do that, we simulated the versus and versus diagrams for a SP composed of 10,000 stars. The errors associated to each magnitude of the simulated stars are extracted from a Gaussian distribution with a dispersion corresponding to the median magnitude error of the RGB stars. Finally, the intrinsic RGB width in the color and in the pseudo-color , indicated respectively as and , have been obtained by subtracting in quadrature the RGB widths of the simulated CMD from observed RGB width ones.
@@ 4.2 Multiple stellar populations in the chromosome map
Similarly to Fig. 4, the red dashed line in the ChM of Terzan 8 and Arp 2 separates bona-fide 1P and 2P stars. As far as Ruprecht 106 and Terzan 7 are concerned, our procedure results in a homogeneous composition of 1P stars only. Both Terzan 8 and Arp 2 show a relatively high fraction of 1P stars compared to most of the studied Galactic GCs. In particular, in Terzan 8 of the cluster RGB stars belong to the 1P, thus making it one of the Galactic GCs with the largest fractions of 1P stars. The 1P predominance in this cluster, identified as a 1P-dominated globulars by Caloi & D’Antona (2011) based on the analysis of its HB morphology, aligns partially with the spectroscopic findings of Carretta et al. (2014). Carretta et al. estimated the 2P fraction in Terzan 8 to be approximately 6%, with only one out of the 16 analyzed stars showing Na abundance consistent with 2P stars. The discrepancy between this study’s estimated 2P fraction () and the value reported by Carretta et al. (2014) likely arises from the smaller sample size used in the latter analysis. Finally, in the case of Arp 2, 1P stars make up a fraction slightly larger than 50% of the RGB cluster stars.
@@ 5 Metallicity variations among first-population stars
In particular, we note that, although Ruprecht 106 does not show any vertical spread, so no presence of 2P stars, the horizontal extension of its ChM covers a interval of mag, significantly wider than that of the photometric errors. Such a phenomenon is even more pronounced in the case of Terzan 8, where the extension of the 1P candidate stars is mag, while in the case of Arp 2 1P and 2P stars seem to span a comparable interval along the x-axis.
Recent works have shown that the horizontal width of 1P stars in the ChMs is a widespread phenomenon among Galactic GCs and demonstrated that 1P stars are not chemically homogeneous (Milone et al., 2015; D’Antona et al., 2016; Milone et al., 2018b). Papers based on both spectroscopy and photometry have shown that the range of 1P stars is mostly due to metallicity variations (e.g. Marino et al., 2019a; Legnardi et al., 2022; Marino et al., 2023). It is important to emphasize that, while additional spectroscopic analyses are needed to confirm the presence of a metallicity spread in every GC with extended 1P sequence, alternative explanations have been proposed to account for the observed elongation of 1P stars. These alternatives include variations in helium content and the presence of unresolved interacting binaries among evolved stars, such as blue stragglers and red giants. However, the observational properties of these stars in the CMD cannot fully account for the spread observed among 1P stars in the ChM (Marino et al., 2019b, 2023; Martins et al., 2020; Kamann et al., 2020; Lardo et al., 2022; Legnardi et al., 2022).
@@ 6 Summary and conclusions
We derived the versus ChMs to search for MPs with different light-element abundances. We find that Arp 2 and Terzan 8 host MPs, whereas Ruprecht 106 and Terzan 7 are SP clusters. This is the first evidence of simple stellar population GCs based on the ChM, and corroborates previous conclusions based on spectroscopy of nine stars of Ruprecht 106 (Villanova et al., 2013), photometry in the U, B, and I bands of RGB stars in both Ruprecht 106 and Terzan 7 (Dotter et al., 2018; Lagioia et al., 2019a), and the HB morphology of Ruprecht 106 (Milone et al., 2014).
@@
It is widely accepted that the pseudo-color width of the RGB correlates with the average cluster metallicity. In GCs with MPs, the RGB width mostly depends on the maximum internal variation in the abundances of CNO. This color index is primarily sensitive to the nitrogen content of the stellar populations and, to a lesser extent, to variations in helium, carbon, and oxygen (e.g. Milone et al., 2015; Lagioia et al., 2019a; Jang et al., 2021). In addition to light-element spread, star-to-star metallicity variations can also contribute to the RGB width (e.g. Marino et al., 2019b). We find that the observed RGB widths of Arp 2 and Terzan 8 follow the same trend as the bulk of Galactic GCs in the versus [Fe/H] diagram. On the contrary, the RGB width of Ruprecht 106 and Terzan 7 are much narrower than those of GCs with similar [Fe/H] as expected for clusters with homogeneous abundance of light-elements.
@@
Previous works have suggested that a mass threshold could govern the occurrence of MPs in GCs. Based on analysis of the ChMs in 58 Galactic GCs, Milone et al. (2020) noticed that the MP GCs have masses larger than . Arp 2 is the Galactic GC with the smallest initial mass ( ) for which MPs have been detected through the ChMs. The discovery of MPs in this cluster suggests that the proposed mass threshold, if present, should be smaller than . In this context, it is worth noticing that various SP Magellanic Cloud GCs have initial masses of 1 – 3 , which are larger than those of Arp 2 and Terzan 8. This fact could indicate that, in addition to cluster mass, the GC formation environment, affects the occurrence of MPs with metal-poor GCs having a lower initial-mass threshold for the formation of 2P stars (Krause et al., 2016; Huang et al., 2024). In this regard, a possible interpretation for the non occurrence of MPs in some GCs is that the emergence of star clusters with MPs may also depend on the environment (Krause et al., 2016; Huang et al., 2024), though it is not obvious how the environment could affect the process. An alternative explanation for the lack of MPs is that this phenomenon would only occur in stars less massive than only (Bastian & Lardo, 2018). In such a case case, the reason for the lack of detection of 2P stars in some Magellanic Cloud GCs is that MP studies in Magellanic Cloud GCs younger than Gyr are exclusively based on RGB stars.
@@
The 1P stars of the multiple-population GCs Arp 2 and Terzan 8 exhibit intrinsic color spreads, which likely indicates the presence of star-to-star [Fe/H] variations of and 0.3 dex, respectively. A recent study of 55 Galactic GCs with MPs by Legnardi et al. (2022) has shown that the 1P stars of virtually all clusters exhibit extended sequences in the ChM. These observations are consistent with inhomogeneous [Fe/H] abundance, ranging from less than 0.05 dex to dex. Hence, the results on Arp 2 and Terzan 8 corroborate the evidence that metallicity variations among 1P stars are common features of GCs.
@@
We found that the SP GCs Ruprecht 106 and Terzan 7 also exhibit extended 1P sequences in the ChM, closely resembling those observed in 1P stars of MP GCs. This finding, combined with evidence of star-to-star [Fe/H] variations in the open cluster NGC 6791 and the Gyr-old star cluster NGC 1783 (Legnardi et al., 2024), indicates that this phenomenon is not exclusive to MP GCs but also occur in SP GCs, open clusters, and intermediate-age massive Magellanic Cloud clusters. Although high-precision spectroscopic observations are required to definitively determine the chemical composition of these systems, star-to-star metallicity variations remain the most plausible explanation for the extended sequences observed in their ChMs.